\begin{thebibliography}{16}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Liu et~al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis, Zettlemoyer, and Stoyanov]{roberta}
Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov.
\newblock Roberta: A robustly optimized bert pretraining approach.
\newblock \emph{ArXiv}, abs/1907.11692, 2019.
\newblock URL \url{https://api.semanticscholar.org/CorpusID:198953378}.

\bibitem[Schweter(2020)]{berturk}
Stefan Schweter.
\newblock Berturk - bert models for turkish, apr 2020.
\newblock URL \url{https://doi.org/10.5281/zenodo.3770924}.

\bibitem[Devlin et~al.(2019)Devlin, Chang, Lee, and Toutanova]{bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova.
\newblock Bert: Pre-training of deep bidirectional transformers for language understanding.
\newblock In \emph{North American Chapter of the Association for Computational Linguistics}, 2019.
\newblock URL \url{https://api.semanticscholar.org/CorpusID:52967399}.

\bibitem[Clark et~al.(2020)Clark, Luong, Le, and Manning]{electra}
Kevin Clark, Minh-Thang Luong, Quoc~V. Le, and Christopher~D. Manning.
\newblock Electra: Pre-training text encoders as discriminators rather than generators.
\newblock In \emph{International Conference on Learning Representations}, 2020.
\newblock URL \url{https://openreview.net/forum?id=r1xMH1BtvB}.

\bibitem[Kudo and Richardson(2018)]{sentencepiece_bpe}
Taku Kudo and John Richardson.
\newblock Sentencepiece: A simple and language independent subword tokenizer and detokenizer for neural text processing.
\newblock \emph{arXiv preprint arXiv:1808.06226}, 2018.

\bibitem[Sennrich et~al.(2015)Sennrich, Haddow, and Birch]{bpe}
Rico Sennrich, Barry Haddow, and Alexandra Birch.
\newblock Neural machine translation of rare words with subword units.
\newblock \emph{ArXiv}, abs/1508.07909, 2015.
\newblock URL \url{https://api.semanticscholar.org/CorpusID:1114678}.

\bibitem[Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin]{transformer}
Ashish Vaswani, Noam~M. Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan~N. Gomez, Lukasz Kaiser, and Illia Polosukhin.
\newblock Attention is all you need.
\newblock In \emph{Neural Information Processing Systems}, 2017.
\newblock URL \url{https://api.semanticscholar.org/CorpusID:13756489}.

\bibitem[Kingma and Ba(2014)]{adam}
Diederik~P. Kingma and Jimmy Ba.
\newblock Adam: A method for stochastic optimization.
\newblock \emph{CoRR}, abs/1412.6980, 2014.
\newblock URL \url{https://api.semanticscholar.org/CorpusID:6628106}.

\bibitem[Ott et~al.(2019)Ott, Edunov, Baevski, Fan, Gross, Ng, Grangier, and Auli]{fairseq}
Myle Ott, Sergey Edunov, Alexei Baevski, Angela Fan, Sam Gross, Nathan Ng, David Grangier, and Michael Auli.
\newblock fairseq: A fast, extensible toolkit for sequence modeling.
\newblock In \emph{North American Chapter of the Association for Computational Linguistics}, 2019.
\newblock URL \url{https://api.semanticscholar.org/CorpusID:91184134}.

\bibitem[Ortiz~Su{\'a}rez et~al.(2020)Ortiz~Su{\'a}rez, Romary, and Sagot]{turkish_oscar}
Pedro~Javier Ortiz~Su{\'a}rez, Laurent Romary, and Beno{\^\i}t Sagot.
\newblock A monolingual approach to contextualized word embeddings for mid-resource languages.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics}, pages 1703--1714, Online, jul 2020. Association for Computational Linguistics.
\newblock URL \url{https://www.aclweb.org/anthology/2020.acl-main.156}.

\bibitem[Raffel et~al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena, Zhou, Li, and Liu]{c4_turkish}
Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter~J. Liu.
\newblock Exploring the limits of transfer learning with a unified text-to-text transformer.
\newblock \emph{Journal of Machine Learning Research}, 21\penalty0 (140):\penalty0 1--67, 2020.
\newblock URL \url{http://jmlr.org/papers/v21/20-074.html}.

\bibitem[Bird et~al.(2009)Bird, Loper, and Klein]{nltk}
Steven Bird, Edward Loper, and Ewan Klein.
\newblock \emph{Natural Language Processing with Python}.
\newblock O'Reilly Media Inc., 2009.

\bibitem[Mar{\c{s}}an et~al.(2022)Mar{\c{s}}an, Akkurt, {\c{S}}en, G{\"u}rb{\"u}z, G{\"u}ng{\"o}r, {\"O}zate{\c{s}}, {\"U}sk{\"u}darl{\i}, {\"O}zg{\"u}r, G{\"u}ng{\"o}r, and {\"O}zt{\"u}rk]{boun_dataset}
B{\"u}{\c{s}}ra Mar{\c{s}}an, Salih~Furkan Akkurt, Muhammet {\c{S}}en, Merve G{\"u}rb{\"u}z, Onur G{\"u}ng{\"o}r, {\c{S}}aziye~Bet{\"u}l {\"O}zate{\c{s}}, Suzan {\"U}sk{\"u}darl{\i}, Arzucan {\"O}zg{\"u}r, Tunga G{\"u}ng{\"o}r, and Balk{\i}z {\"O}zt{\"u}rk.
\newblock Enhancements to the boun treebank reflecting the agglutinative nature of turkish.
\newblock \emph{arXiv preprint arXiv:2207.11782}, 2022.

\bibitem[Sulubacak and Eryiğit(2018)]{imst1}
Umut Sulubacak and Gülşen Eryiğit.
\newblock Implementing universal dependency, morphology and multiword expression annotation standards for turkish language processing.
\newblock \emph{Turkish Journal of Electrical Engineering \& Computer Sciences}, pages 1--23, 5 2018.
\newblock \doi{10.3906/elk-1706-81}.

\bibitem[Sulubacak et~al.(2016)Sulubacak, Gökırmak, Tyers, Çöltekin, Nivre, and Eryiğit]{imst2}
Umut Sulubacak, Memduh Gökırmak, Francis Tyers, Çağrı Çöltekin, Joakim Nivre, and Gülşen Eryiğit.
\newblock Universal dependencies for turkish.
\newblock In \emph{Proceedings of COLING 2016, the 26th International Conference on Computational Linguistics}, Osaka, Japan, 12 2016.

\bibitem[Pan et~al.(2017)Pan, Zhang, May, Nothman, Knight, and Ji]{xtreme}
Xiaoman Pan, Boliang Zhang, Jonathan May, Joel Nothman, Kevin Knight, and Heng Ji.
\newblock Cross-lingual name tagging and linking for 282 languages.
\newblock In Regina Barzilay and Min-Yen Kan, editors, \emph{Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 1946--1958, Vancouver, Canada, jul 2017. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/P17-1178}.
\newblock URL \url{https://aclanthology.org/P17-1178}.

\end{thebibliography}

\documentclass{article}

\usepackage{float}

\usepackage[preprint]{neurips_2023}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{xcolor}         

\title{RoBERTurk: Adjusting RoBERTa for Turkish}

\author{%
  Nuri Tas \\
  Department of Physics\\
  Bogazici University\\
  \texttt{nuri.tas@boun.edu.tr} \\
}

\begin{document}
\maketitle

\begin{abstract}
  We pretrain RoBERTa (\cite{roberta}) on a Turkish corpora using BPE tokenizer. Our model outperforms BERTurk family models  (\cite{berturk}) on the BOUN dataset for the POS task while resulting in underperformance on the IMST dataset for the same task and achieving competitive scores on the Turkish split of the XTREME dataset for the NER task - all while being pretrained on smaller data than its competitors. We release our pretrained model and tokenizer.\footnote{\url{https://huggingface.co/Nuri-Tas/roberturk-base}}
\end{abstract}

\section{Introduction}

Language models such as BERT (\cite{bert}), ELECTRA (\cite{electra}), and RoBERTa (\cite{roberta}) established significant results. However, the careful evaluation of the architectures of these models for the morphologically rich nature of Turkish still needs further investigation. Different characteristics of Turkish, such as flexible word order and agglutinative process, may hinder the performance of contemporary language models, especially in the context of masking algorithms. 

We present a replication of RoBERTa (\cite{roberta}) using Sentencepiece BPE tokenizer (\cite{sentencepiece_bpe}). Our model either outperforms models trained on various Turkish corpora by BERTurk (\cite{berturk}) on the part of speech (POS) tagging task despite being pretrained on a smaller dataset.

\section{Background}
\label{gen_inst}

We briefly review the architecture of RoBERTa (\cite{roberta}) in this section.

The inputs to RoBERTa are tokenized using Byte-Pair Encoding (BPE) (\cite{bpe}) with $50K$ vocabulary size without any preprocessing steps. Tokens are additionally appended with $[BOS]$ and $[EOS]$ special tokens, which denote the beginning and end of a sentence, respectively. Sentences are contiguous text and do not have to be linguistic sentences. 

A random sample of input sequences is then masked with another special token $[MASK]$. Unlike BERT, however, RoBERTa implements a dynamic masking algorithm where new mask patterns are attained at each iteration. Inputs finally go through the transformer model (\cite{transformer}) with $L$ layers using $A$ self-attention
heads and $H$ hidden units without any labels. The pretraining objective is cross-entropy loss of predicting the masked tokens. RoBERTa also removes the next sentence prediction objective during pretraining.

\paragraph{Optimization} RoBERTa uses Adam optimizer (\cite{adam})  with $\epsilon = 1e-6$, $\beta_1 = 0.9$, and $\beta_2 = 0.98$. The learning rate for RoBERTa$_{BASE}$ 
is warmed up to the peak value of $6e-4$ for the first $24K$ updates and linearly decayed to 0. The model is pretrained for maximum $500K$ updates only with sequences of at most $T=512$ length. Note that optimization parameters for fine-tuning on downstream tasks may differ.

\section{Our Setup}
\label{headings}

\paragraph{Implementation}

We use {\fontsize{9}{12}\fontfamily{ptm}\selectfont 
FAIRSEQ%
} (\cite{fairseq}) to pretrain RoBERTa with mixed precision arithmetic.
The model is warmed to the peak value of $1e-5$ for the first $10K$ steps and pretrained for a total of $600K$ steps with mini batches containing 256 samples of maximum length $T=256$.

For tokenization, we use sentencepiece (\cite{sentencepiece_bpe}) library and train BPE on randomly sampled 30M sentences from the training data, which contained around 90M sentences. 

\begin{table}[H]
\caption{Hyperparameters for pretraining RoBERTurk}
\label{hyperparam-pretraining}
    \centering
    \begin{tabular}{lc}
                \toprule
        Hyperparam & RoBERTurk \\
        \midrule
Number of Layers & 12 \\
Hidden size & 1024 \\
FFN inner hidden size & 3072 \\
Attention heads & 12 \\
Attention head size & 64 \\
Dropout & 0.1 \\
Attention Dropout & 0.1 \\
Warmup Steps & 10k \\
Peak Learning Rate & 1e-5 \\
Batch Size & 256 \\
Weight Decay & 0.01 \\
Max Steps & 600k \\
Learning Rate Decay & Linear \\
Adam $\epsilon$ & 1e-6 \\
Adam $\beta_1$ & 0.9 \\
Adam $\beta_2$ & 0.98 \\
        \bottomrule
    \end{tabular}
\end{table}

\subsection{Data}

We make use of two datasets, having a total of 5B tokens and 28GB data size:

• OSCAR (\cite{turkish_oscar}) Turkish split, which is deduplicated.\footnote{The dataset is on  
\href{https://huggingface.co/datasets/oscar/viewer/unshuffled\_deduplicated\_tr}{https://huggingface.co/datasets/oscar/viewer/unshuffled\_deduplicated\_tr}} (27GB).

• The first 12 files of the Turkish split of the processed version of the C4 dataset (\cite{c4_turkish}).\footnote{The dataset is on 
\href{https://huggingface.co/datasets/allenai/c4}{https://huggingface.co/datasets/allenai/c4}}
 Sentences are extracted using the {\fontsize{10}{12}\fontfamily{ptm}\selectfont 
nltk%
} library (\cite{nltk}) in the same preprocessing way by BERTurk (\cite{berturk}). (1GB).

We also note the pretraining data size for the competitor models in Table~\ref{data-sizes}. 

\begin{table}[H]
\caption{Model Pretraining data size}
\label{data-sizes}
    \centering
    \begin{tabular}{lc}
                \toprule
        Model Name & Pretraining data size \\
        \midrule
BERTurk (cased, 128k)    &  35GB \\
BERTurk (cased, 32k)     &  35GB \\
BERTurk (uncased, 128k)  &  242GB \\
BERTurk (uncased, 32k)   &  242GB \\
ConvBERTurk &  35GB \\
ConvBERTurk mC4 (cased)  &  35GB \\
ConvBERTurk mC4 (uncased) &  35GB \\
DistilBERTurk &  35GB \\
ELECTRA BaseException( &  35GB \\
ELECTRA Base mC4 (cased) &  35GB \\
ELECTRA Base mC4 (uncased) &  242GB \\
ELECTRA Small &  242GB \\
RoBERTurk & 28GB \\
        \bottomrule
    \end{tabular}
\end{table}

\subsection{Evaluation}

We finetune the model for part of speech (POS) tagging on BOUN (\cite{boun_dataset}) and IMST (\cite{imst1} and \cite{imst2}) datasets and for named entity recognition (NER) on XTREME (\cite{xtreme}) Turkish split.  BOUN and IMST datasets are annotated in the Universal Dependencies style, and the task is to classify the corresponding label for each word. Similarly, NER task is to classify named entities for each word.

We present the finetuning results in Table~\ref{boun-results}. The results are the average over five runs.  For BOUN and IMST datasets, accuracy is reported, whereas F1 scores are given for the XTREME dataset. While our model outperformed BERTurk family models on BOUN datasets, it yielded less accuracy than its competitors on IMST. Meanwhile, the model achieved competitive scores on the NER task where BERTurk models remarkably accomplished  around 97\%
accuracy.

The hyperparameters for finetuning are given in  Table~\ref{hyperparam-finetuning}. Unlike RoBERTa, the learning rate is kept the same after the warmup process.  

\begin{table}[H]
\caption{Model performance on finetuning tasks}
\label{boun-results}
    \centering
    \begin{tabular}{lccc}
                \toprule
        Model & BOUN & IMST & XTREME\\
        \midrule
        BERTurk (cased, 128k)    &  91.016 ± 0.60  & 96.846 ± 0.42  & 93.8960 ± 0.16 \\
        BERTurk (cased, 32k)     &  91.490 ± 0.10 & 97.096 ± 0.07 & 93.4706 ± 0.09 \\
        BERTurk (uncased, 128k)  &  91.286 ± 0.09 & 97.060 ± 0.07 & 93.4686 ± 0.08 \\
        BERTurk (uncased, 32k)   &  91.544 ± 0.09 & 97.088 ± 0.05 & 92.9086 ± 0.14 \\
        ConvBERTurk &  91.524 ± 0.07 & 97.346 ± 0.07 & \textbf{93.9286 ± 0.07} \\
        ConvBERTurk mC4 (cased)  &  91.724 ± 0.07 & \textbf{97.426 ± 0.03} & 93.6426 ± 0.15 \\
        ConvBERTurk mC4 (uncased) &  91.484 ± 0.12 & 97.338 ± 0.08 & 93.6206 ± 0.13 \\
        DistilBERTurk &  91.044 ± 0.09 & 96.560 ± 0.05 & 91.5966 ± 0.06 \\
        ELECTRA Base &  91.534 ± 0.11 & 97.232 ± 0.09 & 93.4826 ± 0.17 \\
        ELECTRA Base mC4 (cased) &  91.746 ± 0.11 & 97.380 ± 0.05 & 93.6066 ± 0.12 \\
        ELECTRA Base mC4 (uncased) &  91.178 ± 0.15 & 97.210 ± 0.11 & 92.8606 ± 0.36 \\
        ELECTRA Small &  90.850 ± 0.12 & 95.578 ± 0.10 & 90.8306 ± 0.09 \\
        XLM-R (base)     &  91.862 ± 0.16 & 96.492 ± 0.06 & 92.9586 ± 0.14 \\
        mBERT (cased) &  91.492 ± 0.11 & 95.754 ± 0.05 & 93.2306 ± 0.07 \\
        RoBERTurk & \textbf{91.942 ± 0.13}  & 93.812 ± 0.15 & 93.4052 ± 0.16 \\
        \bottomrule
    \end{tabular}
\end{table}

\begin{table}[H]
\caption{Hyperparameters for finetuning RoBERTurk}
\label{hyperparam-finetuning}
    \centering
    \begin{tabular}{lccc}
                \toprule
        Hyperparam & BOUN & IMST & EXTREME \\
        \midrule
Learning Rate & 5e-5 & 1e--5 & 1e-5 \\
Batch Size & 16 & 16 & 16 \\
Weight Decay & 0.06  & 0.06 & 0.1 \\
Max Epochs & 10 &  10 & 10 \\
Warmup Ratio & 0.1 &  0.06 & 0.1 \\
        \bottomrule
    \end{tabular}
\end{table}

\section{Conclusion}

We pretrain RoBERTa on a Turkish corpus and evaluate our model on three different datasets against BERTurk models. Despite being trained on a larger dataset than its competitors, our model achieves competitive scores on BOUN and XTREME datasets but underperforms on IMST.

We release the pretrained model and tokenizer at \url{https://huggingface.co/Nuri-Tas/roberturk-base}.

\section*{Acknowledgements}

We are immensely grateful to Arkadas Ozakin for providing us access to the GPUs at the Kandilli Research Institute and for his helpful advice. We also thank Onur Gungor for pointing out different studies and for his valuable feedback.

\bibliography{references}


\end{document}